### v1

Abstract
A three-term Polak-Ribière-Polyak conjugate gradient method is first proposed in this paper for solving vector optimization problems. This method can autonomously generate descent directions independent of line search procedures, while retaining the conjugate property. The global convergence of the proposed scheme is established by the generalized Wolfe line search procedure without self-adjusting strategies, regular restarts and convexity assumptions. Finally, numerical experiments illustrating the practical performance of this method are presented.
Keywords: Vector optimization; Pareto optimality; Three-term conjugate gradient method; Polak-Ribière-Polyak conjugate gradient method; Global convergence.
2020 MSC: 90C29, 90C52.
1 Introduction
Nonlinear conjugate gradient (CG) methods, known for their fast convergence and low memory requirements, represent a crucial class of first-order algorithms for solving the unconstrained optimization problem
$\min F(x),\qquad x\in\mathbf{R}^{n},$
where $F:\mathbf{R}^{n}\to\mathbf{R}$ is continuously differentiable. A general CG algorithm generates the iterates sequence by the rule
$x^{k+1}=x^{k}+\alpha_{k}d^{k},$
where $\alpha_{k}>0$ is the step size, and $d^{k}$ is the search direction given by
$d^{k}=\left\{\begin{array}[]{l}-\nabla F(x^{k})\qquad\text{for}\ k=0,\\ -\nabla F(x^{k})+\beta_{k}d^{k-1}\qquad\text{for}\ k\geq 1.\end{array}\right.$
Here, $\beta_{k}$ is called the conjugate parameters. Different values of the conjugate parameter $\beta_{k}$ correspond to different CG methods. Some famous CG methods are listed as follows:
[…]
Usually, the conjugate parameter $\beta_{k}$ should ensure that the search direction $d^{k}$ is a descent direction, i.e., $\langle\nabla F(x^{k}),d^{k}\rangle<0$, for all $k\geq 0$. In some convergence analyses, it is necessary to make $d^{k}$ satisfy the sufficient descent condition, namely,
$\langle\nabla F(x^{k}),d^{k}\rangle\leq h\|\nabla F(x^{k})\|^{2},$
for some $h>0$ and all $k\geq 0$.
It is well known that the CG method with the PRP conjugate parameter performs better than other CG methods in practical applications. The global convergence of the PRP-CG method was established in [33] by using the exact line search procedure under the strong convexity hypothesis of the objective function. However, when the objective function is nonconvex, the exact line search can lead to a nonconvergent sequence based on the PRP or HS methods (see [34]). In [17], Gilbert and Nocedal showed the global convergence of the PRP-CG method under the condition that the conjugate parameter is restricted to be nonnegative, i.e., using the truncated strategy, and the line search satisfies the sufficient descent condition and the standard Wolfe conditions. Dai et al. claimed in [7] that the positiveness constraint of the conjugate parameter cannot be relaxed for the PRP method.
On the other hand, some variations of the CG method, known as the three-term CG method, have been widely studied (see [1, 5, 30, 37, 38, 39]). Zhang et al. proposed a three-term FR-CG method [37], a three-term PRP-CG method [38] given as
$d^{k}=-\nabla F(x^{k})+\beta^{PRP}_{k}d^{k-1}-\frac{\langle\nabla F(x^{k}),d^{k-1}\rangle}{\|\nabla F(x^{k-1})\|^{2}}(\nabla F(x^{k})-\nabla F(x^{k-1})),$
and a three-term HS-CG method [39]. Narushima et al. in [30] gave a general form of a class of three-term CG methods. An attractive property of these three-term CG methods is that the search direction generated by these methods satisfies the sufficient descent condition regardless of line search procedures.
Vector optimization arises from the practical demands of multi-objective decision-making problems, and its theoretical framework is widely applied in the fields of management science, space exploration, engineering design, resource allocation and statistics (see [2, 9, 16, 20, 26, 36]). The unconstrained vector optimization problem, which aims to minimize or maximize a vector-valued function relative to a suitable partial order induced by a cone in the image space, is usually defined in the form
$\min_{E}\leavevmode\nobreak\ \mathcal{F}(x),\qquad x\in\mathbf{R}^{n},$
where $\mathcal{F}:\mathbf{R}^{n}\to\mathbf{R}^{m}$ is a continuously differentiable function, $E\subset\mathbf{R}^{m}$ is a closed, convex and pointed cone with nonempty interior. Here, the concept of optimality has to be replaced by the concept of Pareto-optimality or efficiency (see Section 2 below).
Recently, some traditional first-order methods and second-order methods such as the steepest descent method, the subgradient method, the projected gradient method and Newton-type methods have been extended for vector optimization (see [4, 6, 10, 11, 12]). It is well-known that the CG method is superior to the steepest descent method. Unlike the steepest descent method, which often exhibits the sawtooth phenomenon, the CG method achieves faster convergence by using conjugate line search directions, avoiding inefficient repeated searches. Compared to Newton-type methods, the CG method is computationally simpler as it does not require the computation of the Hessian matrix of the objective function or its inverse. Furthermore, the CG method can be compared with Newton-type methods in terms of the convergence rate. Hence, it is necessary to design CG methods for solving vector optimization problems.
Very recently, the extensions of the CG methods FR, CD, DY, PRP and HS for vector optimization were first proposed in [29]. Subsequently, the Hager-Zhang CG method [19], the Liu-Storey CG method [18], the Dai-Liao CG method [24] and the spectral CG method [21] for vector optimization were established. It is important to note that in these works, the extensions of the CG method for vector optimization may not always yield descent directions, even when employing the exact or Wolfe-type line search procedure. This phenomenon is prevalent in vector optimization, as illustrated by the examples provided in [19, 21, 24]. To ensure the descent property of search directions in these studies, the redefinition of conjugate parameters is required, or sufficiently accurate line searches are needed, which means that the descent property of search directions depends on the line search method. Therefore, it is valuable to explore CG methods where the generated search direction inherently satisfies the descent property without relying on the line search procedure.
Motivated by the above works, we first propose a three-term PRP-CG method for vector optimization. The main advantage of this method is that the search direction satisfies the sufficient descent property without redefining the conjugate parameter and depending on the line search procedure, while preserving the conjugate property. It is emphasized that this work can be seen as the first generalization of the traditional three-term CG method to vector optimization, which is the main novelty of this paper. In addition, compared with the stronger line search procedures used in [19, 29], we use the generalized Wolfe line search procedure, without the self-adjusting strategy (e.g., [19, 21]), regular restart (e.g., [18]) and convexity assumption (e.g., [24]), establishing the global convergence of the method. The performance of numerical experiments indicates that the proposed scheme has the excellent characteristics of traditional three-term CG methods for scalar optimization.
The rest of this paper is organized as follows. In Section 2, some notations, concepts and preliminary results for vector optimization are presented. In Section 3, a vector extension of the three-term PRP-CG method is proposed, where the search direction naturally satisfies the sufficient descent property independent of any line search procedure. In Section 4, the global convergence of a three-term PRP-CG algorithm for vector optimization is established using the generalized Wolfe line search method. In Section 5, numerical experiments are provided to show the applicability of the scheme. Final remarks and future directions are discussed in Section 6.
2 Preliminaries
Let $E\subset\mathbf{R}^{m}$ be a closed, convex and pointed cone with nonempty interior. Define the partial order $\preceq_{E}$ (resp. $\prec_{E}$) in $\mathbf{R}^{m}$ by
$u\preceq_{E}v\Leftrightarrow v-u\in E\leavevmode\nobreak\ (\text{resp}.\leavevmode\nobreak\ u\prec_{E}v\Leftrightarrow v-u\in\text{int}(E)).$
In this work, we mainly consider the unconstrained vector optimization problem
$\min_{E}\leavevmode\nobreak\ \mathcal{F}(x),\qquad x\in\mathbf{R}^{n},$ (2.1)
where $\mathcal{F}:\mathbf{R}^{n}\to\mathbf{R}^{m}$ is a continuously differentiable function. We call a point $x^{*}\in\mathbf{R}^{n}$ $E$-Pareto-optimal or $E$-efficient iff there is no $x\in\mathbf{R}^{n}$ such that $\mathcal{F}(x)\preceq_{E}\mathcal{F}(x^{*})$ and $\mathcal{F}(x)\neq\mathcal{F}(x^{*})$. A necessary condition for the local Pareto-optimality of $x\in\mathbf{R}^{n}$ is that
$-\text{int}(E)\cap\text{Image}(J\mathcal{F}(x))=\emptyset,$ (2.2)
where $J\mathcal{F}(x)$ is the Jacobian of $\mathcal{F}$ at $x$ and Image$(J\mathcal{F}(x))$ is the image of $\mathbf{R}^{n}$ by $J\mathcal{F}(x)$.
A point $x^{*}\in\mathbf{R}^{n}$ is called $E$-critical for $\mathcal{F}$ when it satisfies condition (2.2). Hence, if $x\in\mathbf{R}^{n}$ is not $E$-critical, there exists $y\in\mathbf{R}^{n}$ such that $J\mathcal{F}(x)y\in-\text{int}(E)$, i.e., there exists $\varepsilon>0$ such that
$\mathcal{F}(x+ty)\prec_{E}\mathcal{F}(x)$
for all $0<t<\varepsilon$. Each such $y$ is said to be an $E$-descent direction for $\mathcal{F}$ at $x$.
We will denote by $E^{*}$ the positive polar cone of $E$, which is given by
$E^{*}:=\{v\in\mathbf{R}^{m}:\langle v,z\rangle\geq 0,\ \forall z\in E\}.$
Since $E$ is closed and convex, we have $E=E^{**}$,
$-E=\{v\in\mathbf{R}^{m}:\langle v,z\rangle\leq 0,\ \forall z\in E^{*}\}$
and
$-\text{int}(E)=\{v\in\mathbf{R}^{m}:\langle v,z\rangle<0,\ \forall z\in E^{*}-\{0\}\}.$
Let $V$ be a compact set belonging to $E^{*}-\{0\}$ such that the cone generated by the convex hull of $V$ equal to $E^{*}$, i.e.,
$\text{cone}(\text{co}(V))=E^{*}.$
Using the facts that int$(E)\neq\emptyset$ and $V\subseteq E^{*}-\{0\}$, we obtain $0\notin$ co$(V)$. Hence,
$-E=\{v\in\mathbf{R}^{m}:\langle v,z\rangle\leq 0,\ \forall z\in V\}$
and
$-\text{int}(E)=\{v\in\mathbf{R}^{m}:\langle v,z\rangle<0,\ \forall z\in V\}.$
In classical optimization (multiobjective optimization), $E=E^{*}=\mathbf{R}_{+}(E=E^{*}=\mathbf{R}_{+}^{m})$, and $V$ can be chosen as $\{1\}$ (a canonical basis of $\mathbf{R}^{m}$).
For the general closed, convex and pointed cone $E$ with nonempty interior, the set
$V:=\{v\in E^{*}:\|v\|=1\}$
has the properties of $\text{cone}(\text{co}(V))=E^{*}$ and $0\notin$ co$(V)$. In what follows, $V$ will denote exactly this set.
Define $\Psi:\mathbf{R}^{m}\to\mathbf{R}$ and $\Phi:\mathbf{R}^{n}\times\mathbf{R}^{n}\to\mathbf{R}$ by
$\Psi(y):=\sup\{\langle y,v\rangle:v\in V\}$
and
$\Phi(x,d):=\Psi(J\mathcal{F}(x)d)=\sup\{\langle J\mathcal{F}(x)d,v\rangle:v\in V\}.$
Owing to the compactness of $V$, the function $\Psi$ is well defined, and so is the function $\Phi$. It can be easily checked that
  • $x$ is an $E$-critical point for $\mathcal{F}$ iff $\Phi(x,d)\geq 0$ for all $d\in\mathbf{R}^{n}$;
  • $d$ is an $E$-descent direction for $\mathcal{F}$ at $x$ iff $\Phi(x,d)<0$.
We now present some properties of the function $\Phi$.
Lemma 2.1 ([18]).
Let $\mathcal{F}:\mathbf{R}^{n}\to\mathbf{R}^{m}$ be a continuously differentiable function. Then the following statements hold:
  1. i
    For any $x,y,z\in\mathbf{R}^{n}$ and $\alpha>0$, we have
    $\Phi(x,y+\alpha z)\leq\Phi(x,y)+\alpha\Phi(x,z).$
  2. ii
    The function $\Phi(\cdot,\cdot)$ is continuous.
  3. iii
    If the Jacobian $J\mathcal{F}$ is $L$-Lipschitz continuous, i.e., $\|J\mathcal{F}(x)-J\mathcal{F}(y)\|\leq L\|x-y\|$ for all $x,y\in\mathbf{R}^{n}$, then $\Phi(\cdot,d)$ is Lipschitz continuous with $L\|d\|$.
The steepest direction $\mathbf{d}:\mathbf{R}^{n}\to\mathbf{R}^{n}$ for vector optimization is defined by
$\mathbf{d}(x):=\arg\min\bigg{\{}\Phi(x,d)+\frac{\|d\|^{2}}{2}:d\in\mathbf{R}^{n}\bigg{\}}.$ (2.3)
The optimal value $\Theta:\mathbf{R}^{n}\to\mathbf{R}$ of the subproblem in (2.3) is given by
$\Theta(x):=\Phi(x,\mathbf{d}(x))+\frac{\|\mathbf{d}(x)\|^{2}}{2}.$
Since $\Phi(x,\cdot)$ is a real closed convex function, $\mathbf{d}(x)$ exists and is unique. In multiobjective optimization, we choose $E=\mathbf{R}^{m}_{+}$ and $V$ as a canonical basis of $\mathbf{R}^{m}$, then $\mathbf{d}(x)$ can be obtained by solving the following convex quadratic problem:
$\min\leavevmode\nobreak\ a+\frac{\|d\|^{2}}{2},$ (2.4)
[…]
The following lemma shows that the functions $\mathbf{d}$ and $\Theta$ can be used to characterize the $E$-critical points of problem (2.1).
Lemma 2.2 ([29]).
Let $\mathcal{F}:\mathbf{R}^{n}\to\mathbf{R}^{m}$ be a continuously differentiable function. Then the following statements hold:
  1. i
    If $x$ is $E$-critical, then $\mathbf{d}(x)=0$ and $\Theta(x)=0$.
  2. ii
    If $x$ is not $E$-critical, then $\mathbf{d}(x)\neq 0$ is an $E$-descent direction for $\mathcal{F}$ at $x$, $\Theta(x)<0$ and
    $\Phi(x,\mathbf{d}(x))<-\frac{\|\mathbf{d}(x)\|^{2}}{2}<0.$
  3. iii
    The functions $\mathbf{d}$ and $\Theta$ are continuous.
We now turn attention to some line search methods for vector optimization.
Definition 2.3.
It is said that a step size $\alpha>0$ satisfies
  • the exact line search condition if
    $\Phi(x+\alpha d,d)=0;$ (2.5)
  • the standard Wolfe conditions if
    $\mathcal{F}(x+\alpha d)\preceq\mathcal{F}(x)+\rho\alpha\Phi(x,d)\xi,$
    $\Phi(x+\alpha d,d)\geq\sigma\Phi(x,d);$
  • the strong Wolfe conditions if
    $\mathcal{F}(x+\alpha d)\preceq\mathcal{F}(x)+\rho\alpha\Phi(x,d)\xi,$ (2.6)
    $|\Phi(x+\alpha d,d)|\leq\sigma|\Phi(x,d)|;$
where $10<\rho<\sigma<$ 1 are given constants, $\xi\in E$ is such that for any $v\in V$,
$0<\langle v,\xi\rangle\leq 1$ (2.7)
We now introduce a new vector extension of Wolfe-type conditions, based on the scalar form proposed in [5].
Definition 2.4.
We call that a step size $\alpha>0$ satisfies the generalized Wolfe conditions if
$\mathcal{F}(x+\alpha d)\preceq\mathcal{F}(x)+\rho\alpha\Phi(x,d)\xi,$ (2.8)
$-\lambda\Phi(x,d)\geq\Phi(x+\alpha d,d)\geq\sigma\Phi(x,d),$
where $10<\rho<\sigma<$ 1 and $\lambda\geq 0$ are given constants, $\xi\in E$ is as defined in Definition 2.3.
It is easy to see that $\xi\in E$ always exists. Indeed, it can be chosen as $\xi:=\hat{\xi}/\hat{M}$, where $\hat{\xi}\in\text{int}(E)$ and $\hat{M}:=\max\{\langle\hat{\xi},v\rangle:v\in V\}$. In particular, we can take $\xi=[1,\ldots,1]^{T}\in\mathbf{R}^{m}$ in the multiobjective optimization case.
3 Three term PRP-CG method for vector optimization
In this section, we present a three-term PRP-CG method for vector optimization and describe its characteristics.
The vector extension of the PRP conjugate parameter is defined by
$\beta_{k}^{PRP}:=\frac{-\Phi(x^{k},\mathbf{d}(x^{k}))+\Phi(x^{k-1},\mathbf{d}(x^{k}))}{-\Phi(x^{k-1},\mathbf{d}(x^{k-1}))},$
which is first proposed in [29] for solving problem (2.1) by Lucambio Pérez and Prudente. Under the descent assumptions, they establish the convergence of the PRP-CG method with the recurrence
$x^{k+1}=x^{k}+\alpha_{k}d^{k}\qquad\text{for}\leavevmode\nobreak\ k\geq 0,$
in which the step size $\alpha_{k}$ satisfies the strong Wolfe conditions, and the search direction $d^{k}$ is given by
$d^{k}:=\left\{\begin{array}[]{l}\mathbf{d}(x^{k})\qquad\text{for}\ k=0,\\ \mathbf{d}(x^{k})+\beta^{PRP+}_{k}d^{k-1}\qquad\text{for}\ k\geq 1,\end{array}\right.$ (3.1)
where
$\beta_{k}^{PRP+}:=\max\bigg{\{}0,\frac{-\Phi(x^{k},\mathbf{d}(x^{k}))+\Phi(x^{k-1},\mathbf{d}(x^{k}))}{-\Phi(x^{k-1},\mathbf{d}(x^{k-1}))}\bigg{\}}.$ (3.2)
In order for CG methods to be well-defined, $d^{k}$ should be a descent direction of $\mathcal{F}$ at $x^{k}$ for all $k\geq 0$, which is equivalent to $\Phi(x^{k},d^{k})<0$ in vector optimization. Some of these methods need the more stringent condition
$\Phi(x^{k},d^{k})\leq h\Phi(x^{k},\mathbf{d}(x^{k}))$ (3.3)
for some $h\in(0,1]$ and for all $k\geq 0$. The condition (3.3) is called the sufficient descent condition.
Note that for the PRP-CG method, the search direction generated by (3.1) may fail to satisfy the sufficient descent condition and may not even be a descent direction, as shown in the following example.
Example 3.1.
Consider problem (2.1) where $E=\mathbf{R}_{+}^{2}$, $V$ is the canonical basis of $\mathbf{R}^{2}$, and $\mathcal{F}:\mathbf{R}^{2}\to\mathbf{R}^{2}$ is given by
$\mathcal{F}_{1}(x_{1},x_{2}):=\frac{x_{1}^{2}+\sin x_{2}}{2}\quad\text{and}\quad\mathcal{F}_{2}(x_{1},x_{2}):=\frac{(x_{1}-1)^{2}-(x_{2}-1)^{2}}{2}.$
Choose $x^{0}=(1.5,0.9)^{T}$. By solving problem (2.4), we obtain $\mathbf{d}(x^{0})=(-0.5,-0.1)^{T}$. Hence, $d^{0}=\mathbf{d}(x^{0})=(-0.5,-0.1)^{T}$. It is easy to check that $\alpha_{0}=3.1669$ satisfies the exact line search condition (2.5), then $x^{1}=x^{0}+\alpha_{0}d^{0}=(-0.0835,0.5833)^{T}$. Therefore, $\mathbf{d}(x^{1})=(0.0835,-0.4173)^{T}$,
$\beta_{1}^{PRP+}=\max\bigg{\{}0,\frac{-\Phi(x^{1},\mathbf{d}(x^{1}))+\Phi(x^{0},\mathbf{d}(x^{1}))}{-\Phi(x^{0},\mathbf{d}(x^{0}))}\bigg{\}}=0.6966,$
and $d^{1}=\mathbf{d}(x^{1})+\beta_{1}^{PRP+}d^{0}=(-0.2649,-0.4870)^{T}$. Since $\Phi(x^{1},d^{1})=0.0840>0$, we deduce that $d^{1}$ is not an $E$-descent direction of $\mathcal{F}$ at $x^{1}$.
The example shows that the descent property of the PRP-CG method may fail in some cases or may be dependent the selection of line search methods, even when using the truncation strategy. In [29], the authors provided a sufficient condition on the conjugate parameter to guarantee the descent property. However, how to ensure the descent property of the PRP-CG method remains an open problem in their work. To deal with this problem, we propose a three-term PRP-CG method for vector optimization, where the search direction is defined by
$d^{k}:=\left\{\begin{array}[]{l}\mathbf{d}(x^{k})\qquad\text{for}\ k=0,\\ \mathbf{d}(x^{k})+\beta^{PRP}_{k}d^{k-1}+\beta^{PRP}_{k}\frac{\Phi(x^{k},d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\mathbf{d}(x^{k})\qquad\text{for}\ k\geq 1.\end{array}\right.$ (3.4)
We first claim that if a search direction is defined as in (3.4), then this search direction inherently satisfies the sufficient descent condition without relying on any line search method.
Theorem 3.2.
If $x^{k}$ is not $E$-critical and the search direction $d^{k}$ is defined as in (3.4), then $d^{k}$ satisfies the sufficient descent condition (3.3) with $h=1$.
Proof.
Since $x^{k}$ is not $E$-critical, we obtain $\mathbf{d}(x^{k})\neq 0$ by Lemma 2.2(ii). The condition (3.3) trivially holds for $d^{0}=\mathbf{d}(x^{0})$. For any $k\geq 1$, it can be deduced from (3.1) that
$\Phi(x^{k},d^{k})=\Phi\bigg{(}x^{k},\mathbf{d}(x^{k})+\beta^{PRP}_{k}d^{k-1}+\beta^{PRP}_{k}\frac{\Phi(x^{k},d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\mathbf{d}(x^{k})\bigg{)}.$ (3.5)
We now consider the following cases:
Case 1: $\beta^{PRP}_{k},\Phi(x^{k},d^{k-1})\geq 0$. By the definition of $\Phi$ and the fact $\Phi(x^{k},\mathbf{d}(x^{k}))<0$ obtained from Lemma 2.2(ii), we have
$\Phi(x^{k},-\mathbf{d}(x^{k}))$ $=\sup\{\langle-J\mathcal{F}(x^{k})\mathbf{d}(x^{k}),z\rangle:z\in V\}$ (3.6)
$\geq\langle-J\mathcal{F}(x^{k})\mathbf{d}(x^{k}),z_{0}\rangle$
$=-\langle J\mathcal{F}(x^{k})\mathbf{d}(x^{k}),z_{0}\rangle$
$=-\Phi(x^{k},\mathbf{d}(x^{k}))>0,$
where $z_{0}\in V$ is such that
$\langle J\mathcal{F}(x^{k})\mathbf{d}(x^{k}),z_{0}\rangle=\sup\{\langle J\mathcal{F}(x^{k})\mathbf{d}(x^{k}),z\rangle:z\in V\}=\Phi(x^{k},\mathbf{d}(x^{k})).$
Combining (3.5), Lemma 2.1(i) and (3.6), we have
$\Phi(x^{k},d^{k})$ $\leq\Phi(x^{k},\mathbf{d}(x^{k}))+\beta^{PRP}_{k}\Phi(x^{k},d^{k-1})+\beta^{PRP}_{k}\frac{\Phi(x^{k},d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\Phi(x^{k},\mathbf{d}(x^{k}))$
$\leq\Phi(x^{k},\mathbf{d}(x^{k}))+\beta^{PRP}_{k}\Phi(x^{k},d^{k-1})+\beta^{PRP}_{k}\frac{\Phi(x^{k},d^{k-1})}{-\Phi(x^{k},\mathbf{d}(x^{k}))}\Phi(x^{k},\mathbf{d}(x^{k}))$
$\leq\Phi(x^{k},\mathbf{d}(x^{k})).$
Case 2: $\beta^{PRP}_{k}\geq 0$ and $\Phi(x^{k},d^{k-1})<0$. It follows from (3.5) and Lemma 2.1(i) that
$\Phi(x^{k},d^{k})$ $\leq\Phi(x^{k},\mathbf{d}(x^{k}))+\beta^{PRP}_{k}\Phi(x^{k},d^{k-1})+\beta^{PRP}_{k}\frac{-\Phi(x^{k},d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\Phi(x^{k},-\mathbf{d}(x^{k}))$
$\leq\Phi(x^{k},\mathbf{d}(x^{k})).$
Case 3: $\beta^{PRP}_{k},\Phi(x^{k},d^{k-1})<0$. Using similar arguments in (3.6), we infer that
$\Phi(x^{k},-d^{k-1})>-\Phi(x^{k},d^{k-1})>0,$
which, combined with (3.5), Lemma 2.1(i) and (3.6), implies that
$\Phi(x^{k},d^{k})$ $\leq\Phi(x^{k},\mathbf{d}(x^{k}))-\beta^{PRP}_{k}\Phi(x^{k},-d^{k-1})+\beta^{PRP}_{k}\frac{\Phi(x^{k},d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\Phi(x^{k},\mathbf{d}(x^{k}))$
$\leq\Phi(x^{k},\mathbf{d}(x^{k}))-\beta^{PRP}_{k}\Phi(x^{k},-d^{k-1})+\beta^{PRP}_{k}\frac{\Phi(x^{k},d^{k-1})}{-\Phi(x^{k},\mathbf{d}(x^{k}))}\Phi(x^{k},\mathbf{d}(x^{k}))$
$\leq\Phi(x^{k},\mathbf{d}(x^{k}))-\beta^{PRP}_{k}\Phi(x^{k},-d^{k-1})+\beta^{PRP}_{k}\frac{\Phi(x^{k},-d^{k-1})}{\Phi(x^{k},\mathbf{d}(x^{k}))}\Phi(x^{k},\mathbf{d}(x^{k}))$
$\leq\Phi(x^{k},\mathbf{d}(x^{k})).$
Case 4: $\beta^{PRP}_{k}<0,\Phi(x^{k},d^{k-1})\geq 0$. Employing similar lines in (3.6), we have
$0>\Phi(x^{k},-d^{k-1})>-\Phi(x^{k},d^{k-1}),$
which, along with (3.5) and Lemma 2.1(i), yields that
$\Phi(x^{k},d^{k})$ $\leq\Phi(x^{k},\mathbf{d}(x^{k}))-\beta^{PRP}_{k}\Phi(x^{k},-d^{k-1})+\beta^{PRP}_{k}\frac{-\Phi(x^{k},d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\Phi(x^{k},-\mathbf{d}(x^{k}))$
$\leq\Phi(x^{k},\mathbf{d}(x^{k}))-\beta^{PRP}_{k}\Phi(x^{k},-d^{k-1})+\beta^{PRP}_{k}\frac{\Phi(x^{k},-d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\Phi(x^{k},-\mathbf{d}(x^{k}))$
$\leq\Phi(x^{k},\mathbf{d}(x^{k})).$
Combining the above cases, we obtain $\Phi(x^{k},d^{k})\leq\Phi(x^{k},\mathbf{d}(x^{k}))$, which establishes the theorem.
4 Global convergence
In this section, a three-term PRP-CG algorithm for vector optimization is proposed as follows, which satisfies the sufficient descent property and employs the generalized Wolfe line search procedure without the self-adjusting strategy and regular restart.
  Algorithm TT-PRP: A three-term PRP-CG method for vector optimization
  Step 1. Let $10<\rho<\sigma<$ 1, $\lambda>0$, $\xi\in E$ as in (2.7). Compute $\mathbf{d}(x^{0})$ as in (2.3). If $\mathbf{d}(x^{0})=0$, then stop. Otherwise, set $d^{0}:=\mathbf{d}(x^{0})$ and $k\leftarrow 1$.
Step 2. Compute a step size $\alpha_{k-1}>0$ such that
$\mathcal{F}(x^{k})\preceq_{E}\mathcal{F}(x^{k-1})+\rho\alpha_{k-1}\Phi(x^{k-1},d^{k-1})\xi,$ (4.1a)
$-\lambda\Phi(x^{k-1},d^{k-1})\geq\Phi(x^{k},d^{k-1})\geq\sigma\Phi(x^{k-1},d^{k-1}),$ (4.1b)
where
$x^{k}:=x^{k-1}+\alpha_{k-1}d^{k-1}.$ (4.2)
Step 3. Compute $\mathbf{d}(x^{k})$ as in (2.3). If $\mathbf{d}(x^{k})=0$, then stop.
Step 4. Define
$d^{k}:=\mathbf{d}(x^{k})+\beta_{k}d^{k-1}+\frac{\beta_{k}\Phi(x^{k},d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\mathbf{d}(x^{k}),$ (4.3)
where
$\beta_{k}:=\max\bigg{\{}0,\frac{-\Phi(x^{k},\mathbf{d}(x^{k}))+\Phi(x^{k-1},\mathbf{d}(x^{k}))}{-\Phi(x^{k-1},\mathbf{d}(x^{k-1}))}\bigg{\}}.$ (4.4)
Step 5. Let $k\leftarrow k+1$, and go to Step 2.
The following general assumptions will be needed:
$(H_{1})$: The objective function $\mathcal{F}$ is continuously differentiable and the set $V$ is finite and its elements have unit norm.
$(H_{2})$: The level set $\Delta:=\{x\in\mathbf{R}^{n}:\mathcal{F}(x)\preceq_{E}\mathcal{F}(x^{0})\}$ is bounded, i.e., there exists $M>0$ such that $\|x\|\leq M$ for all $x\in\Delta$.
$(H_{3})$: The Jacobian $J\mathcal{F}$ is $L$-Lipschitz continuous on an open set $\Omega\supset\Delta$.
We now consider the convergence of Algorithm TT-PRP. In what follows, we assume that the convergence does not occur in finite steps, i.e., $\mathbf{d}(x^{k})\neq 0$ for all $k$.
By Theorem 3.2, we can easily check that Algorithm TT-PRP generates an infinite sequence $\{(x^{k},d^{k})\}$ with $x^{k}$ and $d^{k}$ satisfying (3.3) for all $k\geq 0$.
Note that if $d^{k}$ is an $E$-descent direction of $\mathcal{F}$ at $x^{k}$ for all $k$, then it can be deduced form assumption $(H_{2})$ that the sequence $\{\Phi(x^{k},\mathbf{d}(x^{k}))\}$ is bounded. In fact, in this case, by the boundedness of $\{x^{k}\}$ and the continuity conclusions of Lemma 2.2, there exist $\gamma,\theta>0$ such that
$\|J\mathcal{F}(x^{k})\|\leq\gamma\quad\text{and}\quad\|\mathbf{d}(x^{k})\|\leq\theta$ (4.5)
for all $k$. Hence, using the continuity conclusions of Lemma 2.1, we can obtain the boundedness of $\{\Phi(x^{k},\mathbf{d}(x^{k}))\}$.
Lemma 4.1.
Under assumptions $(H_{1})-(H_{3})$, if Algorithm TT-PRP generates an infinite sequence $\{x^{k},d^{k}\}$, then
$\sum_{k\geq 0}\frac{\Phi^{2}(x^{k},d^{k})}{\|d^{k}\|^{2}}<\infty.$
Proof.
Since $d^{k}$ is an $E$-descent direction of $\mathcal{F}$ at $x^{k}$ for all $k\geq 0$, we have that $d^{k}\neq 0$ and $\Phi^{2}(x^{k},d^{k})/\|d^{k}\|^{2}$ is well defined. Since the step size $\alpha_{k}$ satisfies the generalized Wolfe conditions (LABEL:2*) for all $k\geq 0$, we can use [29, Proposition 3.3] to obtain the desired conclusion. ∎
In what follows, we will claim that $\mathbf{d}(x^{k})$ cannot be bounded away from zero. To do this, we need to make the following contradictory assumption:
$(H_{4})$: There exists $\eta>0$ such that for any $k\geq 0$,
$\|\mathbf{d}(x^{k})\|\geq\eta.$
We present some auxiliary lemmas.
Lemma 4.2.
Assume that assumptions $(H_{1})-(H_{4})$ hold, if Algorithm TT-PRP generates an infinite sequence $\{x^{k},d^{k}\}$, then
$\sum^{\infty}_{k=0}\frac{1}{\|d^{k}\|^{2}}<\infty.$
Proof.
By assumption $(H_{4})$, Lemma 2.2(ii) and Theorem 3.2, for any $k\geq 0$, we have
$\frac{\eta^{4}}{4\|d^{k}\|^{2}}\leq\frac{\|\mathbf{d}(x^{k})\|^{4}}{4\|d^{k}\|^{2}}\leq\frac{\Phi^{2}(x^{k},\mathbf{d}(x^{k}))}{\|d^{k}\|^{2}}\leq\frac{\Phi^{2}(x^{k},d^{k})}{\|d^{k}\|^{2}},$
which, using Lemma 4.1, yields that
$\frac{{\eta}^{4}}{4}\sum^{\infty}_{k=0}\frac{1}{\|d^{k}\|^{2}}<\infty.$
Lemma 4.3.
Under assumptions $(H_{1})-(H_{4})$, if Algorithm TT-PRP generates an infinite sequence $\{x^{k},d^{k}\}$, then
$\sum_{k\geq 1}\|u^{k}-u^{k-1}\|<\infty,$
where $u^{k}:=d^{k}/\|d^{k}\|$.
Proof.
Since (4.3) and the definition of $\Phi$, we have
$d^{k}=\mathbf{d}(x^{k})+\|d^{k-1}\|\bigg{(}\beta_{k}u^{k-1}+\frac{\beta_{k}\Phi(x^{k},u^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\mathbf{d}(x^{k})\bigg{)}.$
From this, we can assume that
$\|d^{k}\|=\mu_{k}(x^{k})+\nu_{k}(x^{k})\|d^{k-1}\|$ (4.6)
for all $k\geq 1$, where $\mu_{k},\nu_{k}:\mathbf{R}^{n}\to\mathbf{R}$. Using the fact $\|u^{k-1}\|=1$, the boundedness of $\{x^{k}\}$, and the continuity conclusions of Lemmas 2.1 and 2.3, we can infer that the sequences $\{\mu_{k}(x^{k})\}$ and $\{\nu_{k}(x^{k})\}$ are bounded. Hence, there exists $C_{\mu}>0$ such that
$|\mu_{k}(x^{k})|\leq C_{\mu}$
for all $k\geq 1$. By the boundedness of $\{\nu_{k}(x^{k})\}$, we obtain that $\liminf_{k\to\infty}\nu_{k}(x^{k})$ exists. Moreover, we can assert that
$\omega:=\liminf_{k\to\infty}\nu_{k}(x^{k})>0.$
Otherwise, it follows from (4.6) that
$\liminf_{k\to\infty}\|d^{k}\|<\liminf_{k\to\infty}\mu_{k}(x^{k})\leq C_{\mu}<\infty,$
which contrary to Lemma 4.2. Based on this, we divide the sequence $\{k\}$ into the two subsequences
$\{\overline{k}:\overline{k}\geq 1,\nu_{\overline{k}}(x^{\overline{k}})\geq\omega\}$
and
$\{\underline{k}:\underline{k}\geq 1,\nu_{\underline{k}}(x^{\underline{k}})<\omega\}.$
By the definition of the limit inferior, the sequence $\{\underline{k}\}$ is finite. According to Lemma 4.2, there exists $N_{0}>0$ such that for any $k\geq N_{0}$,
$\|d^{k}\|>\frac{2C_{\mu}+1}{\omega}.$ (4.7)
Let
$\varpi^{k}:=\frac{1}{\|d^{k}\|}\bigg{(}\mathbf{d}(x^{k})+\frac{\beta_{k}\Phi(x^{k},d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\mathbf{d}(x^{k})\bigg{)}.$
From (4.3), we have
$u^{k}=\varpi^{k}+\beta_{k}\frac{\|d^{k-1}\|}{\|d^{k}\|}u^{k-1}.$ (4.8)
Using the fact $\|u^{k}\|=\|u^{k-1}\|=1$, we deduce from (4.8) that
$\|\varpi^{k}\|=\bigg{\|}u^{k}-\beta_{k}\frac{\|d^{k-1}\|}{\|d^{k}\|}u^{k-1}\bigg{\|}=\bigg{\|}\beta_{k}\frac{\|d^{k-1}\|}{\|d^{k}\|}u^{k}-u^{k-1}\bigg{\|},$
which, combined with $\beta_{k}\geq 0$, gives that
$\|u^{k}-u^{k-1}\|$ $\leq\bigg{(}\beta_{k}\frac{\|d^{k-1}\|}{\|d^{k}\|}+1\bigg{)}\|u^{k}-u^{k-1}\|$ (4.9)
$=\bigg{\|}u^{k}-\beta_{k}\frac{\|d^{k-1}\|}{\|d^{k}\|}u^{k-1}+\beta_{k}\frac{\|d^{k-1}\|}{\|d^{k}\|}u^{k}-u^{k-1}\bigg{\|}$
$\leq\bigg{\|}\beta_{k}\frac{\|d^{k-1}\|}{\|d^{k}\|}u^{k}-u^{k-1}\bigg{\|}+\bigg{\|}u^{k}-\beta_{k}\frac{\|d^{k-1}\|}{\|d^{k}\|}u^{k-1}\bigg{\|}$
$\leq 2\|\varpi^{k}\|.$
By (3.3), (4.6), (4.7), the continuity conclusions of Lemmas 2.1 and 2.2, the boundedness of $\{x^{k}\}$, and Lemmas 4.1 and 4.2, we obtain from (4.9) the following estimate
$\sum^{\infty}_{k=1}\|u^{k}-u^{k-1}\|^{2}$
$\leq\sum^{\infty}_{k=1}\frac{4}{\|d^{k}\|^{2}}\bigg{(}\|\mathbf{d}(x^{k})\|+\frac{\beta_{k}|\Phi(x^{k},d^{k-1})|}{|\Phi(x^{k},-\mathbf{d}(x^{k}))|}\|\mathbf{d}(x^{k})\|\bigg{)}^{2}$
$\leq\sum^{\infty}_{k=1}\frac{8\|\mathbf{d}(x^{k})\|^{2}}{\|d^{k}\|^{2}}+\sum^{\infty}_{k=1}\frac{8\beta^{2}_{k}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{k})\|^{2}}{\Phi^{2}(x^{k},-\mathbf{d}(x^{k}))}\frac{\Phi^{2}(x^{k-1},d^{k-1})}{\|d^{k}\|^{2}}$
$\leq\sum^{\infty}_{k=1}\frac{8\|\mathbf{d}(x^{k})\|^{2}}{\|d^{k}\|^{2}}+\sum_{\overline{k}\leq N_{0}}\frac{8\beta^{2}_{\overline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\overline{k}})\|^{2}}{\Phi^{2}(x^{\overline{k}},-\mathbf{d}(x^{\overline{k}}))}\frac{\Phi^{2}(x^{\overline{k}-1},d^{\overline{k}-1})}{\|d^{\overline{k}}\|^{2}}$
$\quad+\sum_{\overline{k}>N_{0}}\frac{8\beta^{2}_{\overline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\overline{k}})\|^{2}}{\Phi^{2}(x^{\overline{k}},-\mathbf{d}(x^{\overline{k}}))}\frac{\Phi^{2}(x^{\overline{k}-1},d^{\overline{k}-1})}{\|d^{\overline{k}}\|^{2}}$
$\quad+\sum_{\underline{k}}\frac{8\beta^{2}_{\underline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\underline{k}})\|^{2}}{\Phi^{2}(x^{\underline{k}},-\mathbf{d}(x^{\underline{k}}))}\frac{\Phi^{2}(x^{\underline{k}-1},d^{\underline{k}-1})}{\|d^{\underline{k}}\|^{2}}$
$\leq\sum^{\infty}_{k=1}\frac{8\|\mathbf{d}(x^{k})\|^{2}}{\|d^{k}\|^{2}}+\sum_{\overline{k}\leq N_{0}}\frac{8\beta^{2}_{\overline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\overline{k}})\|^{2}}{\Phi^{2}(x^{\overline{k}},-\mathbf{d}(x^{\overline{k}}))}\frac{\Phi^{2}(x^{\overline{k}-1},d^{\overline{k}-1})}{\|d^{\overline{k}}\|^{2}}$
$\quad+\sum_{\overline{k}>N_{0}}\frac{32\beta^{2}_{\overline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\overline{k}})\|^{2}}{\Phi^{2}(x^{\overline{k}},-\mathbf{d}(x^{\overline{k}}))}\frac{\Phi^{2}(x^{\overline{k}-1},d^{\overline{k}-1})}{\omega^{2}\|d^{\overline{k}-1}\|^{2}}$
$\quad+\sum_{\underline{k}}\frac{8\beta^{2}_{\underline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\underline{k}})\|^{2}}{\Phi^{2}(x^{\underline{k}},-\mathbf{d}(x^{\underline{k}}))}\frac{\Phi^{2}(x^{\underline{k}-1},d^{\underline{k}-1})}{\|d^{\underline{k}}\|^{2}}$
<<$\infty$.< ∞ .
Here, we use the finiteness of the sequences $\{\overline{k}\leq N\}$ and $\{\underline{k}\}$, and the proof is complete.
Theorem 4.4.
If assumptions $(H_{1})-(H_{3})$ hold and Algorithm TT-PRP generates an infinite sequence $\{x^{k},d^{k}\}$, then
$\liminf_{k\to\infty}\|\mathbf{d}(x^{k})\|=0.$
Proof.
Suppose, contrary to our claim, that assumption $(H_{4})$ holds. By assumption $(H_{3})$, Lemmas 2.1(iii) and 2.2(ii), and (4.5), it follows from (4.6) that
$|\beta_{k}|$ $\leq\frac{|-\Phi(x^{k},\mathbf{d}(x^{k}))+\Phi(x^{k-1},\mathbf{d}(x^{k}))|}{|\Phi(x^{k-1},\mathbf{d}(x^{k-1}))|}$ (4.10)
$\leq\frac{2L\|\mathbf{d}(x^{k})\|\|x^{k}-x^{k-1}\|}{|\mathbf{d}(x^{k-1})|^{2}}$
$\leq\frac{2L\theta}{{\eta}^{2}}\|x^{k}-x^{k-1}\|.$
Moreover, it can be inferred from Lemma 2.2(ii) and (4.5) that
$|\beta_{k}|\leq\frac{|-\Phi(x^{k},\mathbf{d}(x^{k}))+\Phi(x^{k-1},\mathbf{d}(x^{k}))|}{|\Phi(x^{k-1},\mathbf{d}(x^{k-1}))|}\leq\frac{4\gamma\theta}{\eta^{2}}.$ (4.11)
If $\Phi(x^{k},d^{k-1})\geq 0$, we would have
$|\Phi(x^{k},d^{k-1})|\leq|\Phi(x^{k},d^{k-1})-\Phi(x^{k-1},d^{k-1})|.$
If $\Phi(x^{k},d^{k-1})\leq 0$, we deduce from (4.1b) that
$|\Phi(x^{k},d^{k-1})|\leq\sigma|\Phi(x^{k-1},d^{k-1})|,$
which leads to
$|\Phi(x^{k},d^{k-1})|$ $\leq\frac{\sigma}{1-\sigma}(|\Phi(x^{k-1},d^{k-1})|-|\Phi(x^{k},d^{k-1})|)$
$\leq\frac{\sigma}{1-\sigma}|\Phi(x^{k},d^{k-1})-\Phi(x^{k-1},d^{k-1})|.$
Combining these and using Lemma 2.1(iii), we see that
$|\Phi(x^{k},d^{k-1})|$ $\leq C_{1}|\Phi(x^{k},d^{k-1})-\Phi(x^{k-1},d^{k-1})|$ (4.12)
$\leq C_{1}L\|x^{k}-x^{k-1}\|\|d^{k-1}\|,$
where $C_{1}:\max\{1,\sigma/(1-\sigma)\}$. From (3.6) and Lemma 2.2(ii), we have
$|\Phi(x^{k},-\mathbf{d}(x^{k}))|\geq|\Phi(x^{k},\mathbf{d}(x^{k}))|>\frac{\|\mathbf{d}(x^{k})\|^{2}}{2}\geq\frac{\eta^{2}}{2}.$ (4.13)
Combining (4.5), (4.10), (4.11), (4.12) and (4.13), we infer from (4.3) that
$\|d^{k}\|^{2}$ $\leq\bigg{(}\|\mathbf{d}(x^{k})\|+\beta_{k}\|d^{k-1}\|+\frac{\beta_{k}|\Phi(x^{k},d^{k-1})|\|\mathbf{d}(x^{k})\|}{|\Phi(x^{k},-\mathbf{d}(x^{k}))|}\bigg{)}^{2}$ (4.14)
$\leq\bigg{(}\theta+\frac{2L\theta}{{\eta}^{2}}\|x^{k}-x^{k-1}\|\|d^{k-1}\|+\frac{8C_{1}L\gamma\theta^{2}}{\eta^{4}}\|x^{k}-x^{k-1}\|\|d^{k-1}\|\bigg{)}^{2}$
$\leq 2{\theta}^{2}+2{C}^{2}\|x^{k}-x^{k-1}\|^{2}\|d^{k-1}\|^{2}$
for all $k\geq 1$, where $C:=(2L\theta{\eta}^{2}+8C_{1}L\gamma\theta^{2})/{\eta}^{4}$.
[…]
$\sum_{i\geq m_{0}+1}\|u^{i}-u^{i-1}\|^{2}\leq\frac{1}{4N}.$ (4.15)
By the definition of $u^{k}$, we deduce that for any $m>m_{0}$,
$x^{m}-x^{m_{0}}$ $=\sum^{m}_{i=m_{0}+1}(x^{i}-x^{i-1})=\sum^{m}_{i=m_{0}+1}\|s^{i-1}\|u^{i-1}$
$=\sum^{m}_{i=m_{0}+1}\|s^{i-1}\|u^{m_{0}}+\sum^{m}_{i=m_{0}+1}\|s^{i-1}\|(u^{i-1}-u^{m_{0}}),$
where $s^{i-1}:=x^{i}-x^{i-1}$. From this, we get
$\sum^{m}_{i=m_{0}+1}\|s^{i-1}\|$ $\leq\|x^{m}-x^{m_{0}}\|+\sum^{m}_{i=m_{0}+1}\|s^{i-1}\|\|u^{i-1}-u^{m_{0}}\|$ (4.16)
$\leq 2M+\sum^{m}_{i=m_{0}+1}\|s^{i-1}\|\|u^{i-1}-u^{m_{0}}\|.$
Since (4.15) and the fact that $(\sum^{n}_{i=1}a_{i})^{2}\leq n\sum^{n}_{i=1}a_{i}^{2}$ with $a_{i}\in\mathbf{R}$, it can be inferred that for any $m>m_{0}$ with $m-m_{0}\leq N$,
$\|u^{m}-u^{m_{0}}\|\leq\sum^{m}_{i=m_{0}+1}\|u^{i}-u^{i-1}\|\leq\sqrt{m-m_{0}}\bigg{(}\sum^{m}_{i=m_{0}+1}\|u^{i}-u^{i-1}\|^{2}\bigg{)}^{\frac{1}{2}}\leq\sqrt{N}\bigg{(}\frac{1}{4N}\bigg{)}^{\frac{1}{2}}=\frac{1}{2},$
which, combined with (4.16), means that
$\sum^{m}_{i=m_{0}+1}\|s^{i-1}\|\leq 4M.$ (4.17)
By (4.14) and the definition of $\{s^{i}\}$, we conclude from induction that
$\|d^{m}\|^{2}\leq 2{\theta}^{2}\bigg{(}\sum^{m+1}_{i=m_{0}+1}\prod^{m}_{j=i}2C^{2}\|s^{j-1}\|^{2}\bigg{)}+\|d^{m_{0}}\|^{2}\prod^{m}_{j=m_{0}+1}2{C}^{2}\|s^{j-1}\|^{2}$ (4.18)
for all $m>m_{0}$. Here, if the index range is empty, then the above product is defined to be one. Using the arithmetic-geometric mean inequality and (4.17), we see that
$\prod^{m_{0}+N}_{j=m_{0}+1}2{C}^{2}\|s^{j-1}\|^{2}=\bigg{(}\prod^{m_{0}+N}_{j=m_{0}+1}\sqrt{2}C\|s^{j-1}\|\bigg{)}^{2}$ (4.19)
$\leq\bigg{(}\frac{\sum^{m_{0}+N}_{j=m_{0}+1}\sqrt{2}C\|s^{j-1}\|}{N}\bigg{)}^{2N}\leq\bigg{(}\frac{4\sqrt{2}CM}{N}\bigg{)}^{2N}\leq\frac{1}{2^{N}}$
for all $N\geq\lceil 8CM\rceil$. Combining (4.18) and (4.19), for any $m\geq m_{0}+\lceil 8CM\rceil$, we have
$\|d^{m}\|^{2}$ $\leq 2{\theta}^{2}\bigg{(}\sum^{m_{0}+\lceil 8CM\rceil}_{i=m_{0}+1}\prod^{m_{0}+\lceil 8CM\rceil-1}_{j=i}2C^{2}\|s^{j-1}\|^{2}\bigg{)}$ (4.20)
$\quad+2\theta^{2}\bigg{(}\sum^{m+1}_{i=m_{0}+\lceil 8CM\rceil+1}\prod^{m}_{j=i-\lceil 8CM\rceil}2C^{2}\|s^{j-1}\|^{2}\bigg{)}+\|d^{m_{0}}\|^{2}\prod^{m}_{j=m_{0}+1}2C^{2}\|s^{j-1}\|^{2}$
$\leq 2\theta^{2}\bigg{(}\sum^{m_{0}+\lceil 8CM\rceil}_{i=m_{0}+1}\prod^{m_{0}+\lceil 8CM\rceil-1}_{j=i}2C^{2}\|s^{j-1}\|^{2}\bigg{)}$
$\quad+2\theta^{2}\bigg{(}\sum^{m+1}_{i=m_{0}+\lceil 8CM\rceil+1}\frac{1}{2^{m+\lceil 8CM\rceil+1-i}}\bigg{)}+\frac{\|d^{m_{0}}\|^{2}}{2^{m-m_{0}}}$
$\leq 2\theta^{2}\bigg{(}\sum^{m_{0}+\lceil 8CM\rceil}_{i=m_{0}+1}\prod^{m_{0}+\lceil 8CM\rceil-1}_{j=i}2C^{2}\|s^{j-1}\|^{2}\bigg{)}+4\theta^{2}+\|d^{m_{0}}\|^{2},$
which implies that $d^{m}$ is bounded independently of $m$. This contradicts Lemma 4.3, and the theorem is proved.∎
5 Numerical experiments
In this section, we give some numerical results to illustrate the effectiveness of Algorithm TT-PRP. All codes are written in MATLAB R2023b and run on a PC with CPU 2.50 GHz, 16 GB RAM. The test problems are selected from multi-objective optimization literature and its relevant references. Hence, we assume that $E:=\mathbf{R}_{+}^{m}$, $V$ is the canonical basis of $\mathbf{R}^{m}$, and $\xi:=[1,\ldots,1]^{T}\in\mathbf{R}^{m}$.
The considered methods in the numerical tests are as follows:
  • TT-PRP CG method: The conjugate parameter and the step size are given by Algorithm TT-PRP.
  • TT-PRP1 CG method: The conjugate parameter is defined by (4.3), and the step size satisfies the strong Wolfe conditions (2.6).
  • PRP+ CG method [29]: The conjugate parameter is given by (3.2), and the step size is obtained by a strong Wolfe line search procedure.
  • The steepest descent (SD) method [15]: The step size is governed by the strong Wolfe conditions (2.6).
In our implementation, the parameters in Definitions 2.3 and 2.4 are chosen as $\rho=0.01$, $\sigma=0.2$ and $\lambda=2$, and the maximum number of allowed iterations is set to 3000.
Motivated by [28], we design a line search procedure that identifies step sizes satisfying the generalized Wolfe conditions (LABEL:2*). The algorithm is briefly outlined as follows. Suppose that $d$ is an $E$-descent direction for $\mathcal{F}$ at $x$. From $\Phi(x,d)<0$ and the definition of $\Phi$, we can assert that $d$ is a descent direction for all scalar functions $\mathcal{F}_{i}(x)$ with $i=1,...,m$. Hence, the first step is that finds $\alpha^{i}>0$ satisfying
$\mathcal{F}_{i}(x+\alpha^{i}d)\leq\mathcal{F}_{i}(x)+\rho\alpha^{i}\Phi(x,d),$ (5.1)
$-\lambda\Phi(x,d)\geq\langle\nabla\mathcal{F}_{i}(x+\alpha^{i}d),d\rangle\geq\sigma\Phi(x,d)$
for some index $i\in\{1,\ldots,m\}$. If $\alpha^{i}$ satisfies (5.1) for all $i\in\{1,\ldots,m\}$, then $\alpha^{i}$ obviously satisfies the generalized Wolfe conditions (LABEL:2*). Otherwise, we choose another scalar function $\mathcal{F}_{j}(x)$ with $j\in\{1,\ldots,m\}$, and seek $\alpha^{j}\in(0,\alpha^{i})$ such that (5.1) holds for $\mathcal{F}_{j}(x)$. The step size conforming to the generalized Wolfe conditions (LABEL:2*) can be obtained by repeating the above process.
By Lemma 2.2, we deduce that $\mathbf{d}(x^{k})=0$ iff $\Theta(x^{k})=0$ and $\mathbf{d}(x^{k})\neq 0$ iff $\Theta(x^{k})<0$. From these, we can stop the algorithms at $x^{k}$ declaring convergence whenever
$\Theta(x^{k})\geq-5\times\text{eps}^{\frac{1}{2}},$
where $\Theta(x^{k})=\Phi(x^{k},\mathbf{d}(x^{k}))+\|\mathbf{d}(x^{k})\|^{2}/2$ and $\text{esp}\approx 2.22\times 10^{-16}$ denotes the machine precision. The steepest descent direction $\mathbf{d}(x^{k})$ and the optimal value $\Theta(x^{k})$ can be obtained by using the built-in function quadprog in Matlab to solve subproblem (2.4).
All considered problems are listed in Table 1. The columns “Problem” and “Source” respectively represent the problem name and the corresponding reference. The column “Convex” indicates the convexity of the test problems. The columns “$n$” and “$m$” represent the numbers of variables and objectives for the problems, respectively. The initial points are taken from a box given in the column “$x_{D}$”.
Table 1: Test problems
Problem    Source    $n$    $m$    Convex    $x_{D}$
AP3 [3] $2$ $2$ N $[-2,2]^{n}$
Far1 [25] $2$ $2$ N $[-1,1]^{n}$
FDS-1 [14] $2$ $3$ Y $[-2,2]^{n}$
FDS-2 [14] $100$ $3$ Y $[-2,2]^{n}$
Hil1 [23] $2$ $2$ N $[0,1]^{n}$
Lov3 [27] $2$ $2$ N $[-100,100]^{n}$
Lov4 [27] $2$ $2$ N $[-100,100]^{n}$
MGH26 [32] $4$ $4$ N $[-1,1]^{n}$
MMR5-1 [31] 1000 2 N $[-5,5]^{n}$
MMR5-2 [31] 200 2 N $[-100,100]^{n}$
MOP5 [25] $2$ $3$ N $[-1,1]^{n}$
MOP7 [25] $2$ $3$ Y $[-400,400]^{n}$
SLC2-1 [35] $1000$ $2$ Y $[-10,10]^{n}$
SLC2-2 [35] $200$ $2$ Y $[-100,100]^{n}$
Every problem was solved 100 times with starting points generated from a uniform random distribution within the box specified in the column “$x_{D}$”. To provide an intuitive representation of the numerical results, Table 2 summarizes the performance of each method for the test problems, including the percentage of runs that reached a critical point (%), the median of iterations (it), the median of function evaluations (evalf), and the median of gradient evaluations (evalg).
As observed, the TT-PRP method exhibits robust convergence characteristics for all test problems. For the problems AP3, FDS-1, Lov3, Lov4, MGH26, MOP5 and MOP7, the methods TT-PRP and TT-PRP1 show comparable performance to the PRP+ method. In some small-scale nonconvex problems, such as Far1 and Hil1, the methods TT-PRP and TT-PRP1 exhibit a slight advantage over the PRP+ method. For large-scale problems, such as FDS-2, MMR5-1, MMR5-2, SLC2-1 and SLC2-2, the methods TT-PRP and TT-PRP1 outperform the methods PRP+ and SD. This suggests that both the sufficient descent property and conjugate property are presented for the methods TT-PRP and TT-PRP1, which are superior to the methods PRP+ and SD. Compared with the TT-PRP1 method, the TT-PRP method has a smaller median of iterations in most test problems, and the medians of function evaluations and gradient evaluations are lower when the median of iterations is similar. This indicates that due to the relaxation of the line search, the TT-PRP method can achieve larger step sizes with fewer gradient evaluations and function evaluations than the TT-PRP1 method.
Table 2: Performance of the considered CG algorithms in the test problems.
TT-PRP TT-PRP1 PRP+ SD TT-PRP TT-PRP1 PRP+ SD
AP3 MGH26
% 100.0 100.0 100.0 100.0 % 100.0 100.0 100.0 100.0
it 7.0 7.0 7.0 69.5 it 6.0 6.0 6.0 8.0
evalf 53.0 55.0 55.0 465.5 evalf 34.0 37.0 37.0 60.0
evalg 43.0 46.0 46.0 398.0 evalg 26.0 29.0 29.0 50.0
Far1 MMR5-1
% 100.0 100.0 100.0 100.0 % 100.0 100.0 100.0 100.0
it 33.0 32.5 38.0 70.5 it 134.0 145.0 183.5 214.0
evalf 276.0 308.0 375.5 523.5 evalf 738.0 964.0 1287.0 1475.0
evalg 242.0 273.0 337.5 451.0 evalg 602.0 837.0 1101.0 1258.0
FDS-1 MMR5-2
% 100.0 100.0 100.0 100.0 % 100.0 100.0 100.0 100.0
it 6.0 6.0 6.0 8.0 it 151.0 163.0 220.0 283.0
evalf 52.0 61.0 61.0 82.0 evalf 784.0 960.5 1682.5 2264.5
evalg 44.0 53.0 53.0 72.0 evalg 611.0 795.0 1460.5 1979.5
FDS-2 MOP5
% 100.0 100.0 100.0 100.0 % 100.0 100.0 100.0 100.0
it 130.0 145.5 171.5 221.5 it 3.0 3.0 3.0 3.0
evalf 779.5 1163.5 1363.5 2687.5 evalf 28.0 28.0 28.0 28.0
evalg 911.5 1311.0 1537.0 2464.0 evalg 23.0 23.0 23.0 23.0
Hil1 MOP7
% 100.0 100.0 100.0 100.0 % 100.0 100.0 100.0 100.0
it 8.0 9.0 10.0 16.5 it 8.0 8.0 8.0 27.5
evalf 55.0 72.0 78.0 102.5 evalf 47.0 58.0 58.0 165.0
evalg 45.0 61.0 66.0 84.0 evalg 37.0 48.0 48.0 135.0
Lov3 SLC2-1
% 100.0 100.0 100.0 100.0 % 100.0 100.0 100.0 100.0
it 3.0 3.0 3.0 3.0 it 20.5 22.0 32.5 62.0
evalf 26.0 26.0 26.0 26.0 evalf 301.0 328.5 517.0 928.5
evalg 21.0 21.0 21.0 21.0 evalg 278.5 304.5 482.5 864.5
Lov4 SLC2-2
% 100.0 100.0 100.0 100.0 % 100.0 100.0 100.0 100.0
it 1.0 1.0 1.0 1.0 it 20.0 21.5 33.0 58.0
evalf 6.0 6.0 6.0 6.0 evalf 212.0 241.0 360.5 772.0
evalg 5.0 5.0 5.0 5.0 evalg 189.0 217.5 325.5 712.0
The ability of various algorithms to successfully obtain Pareto frontiers has always attracted great interest in vector optimization problems with finite objectives. To evaluate the effectiveness of the TT-PRP method in generating Pareto frontiers, the value spaces of the bicriteria problems (AP3, Far1 and Hil1) and the three-criteria problems (FDS-1, MOP5 and MOP7) are plotted in Figures 1 and 2 using the TT-PRP method. For these problems, 400 starting points are randomly generated (except for problem Far1, which uses 600 points due to the complexity of the image space) in the box given in the column “$x_{D}$” in Table 1. In the figures, a green point represents a final iteration, while the beginning of a straight segment represents the corresponding starting point. The results in Figures 1 and 2 show the effective estimation of the Pareto frontier for the considered problems using a reasonable number of initial points by the TT-PRP method.
(a) Image set of AP3
(b) AP3
(c) Image set of Far1
(d) Far1
(e) Image set of Hil1
(f) Hil1
Figure 1: Image sets and value spaces generated by the TT-PRP method for problems AP3, Far1 and Hil1.
(a) Image set of FDS-1
(b) FDS-1
(c) Image set of MOP5
(d) MOP5
(e) Image set of MOP7
(f) MOP7
Figure 2: Image sets and value spaces generated by the TT-PRP method for problems FDS-1, MOP5 and MOP7.
6 Final remarks and future directions
We first propose a three-term Polak-Ribière-Polyak conjugate gradient method for solving vector optimization problems, which can be regarded as the direct extension of the traditional three-term conjugate gradient method. It is worth mentioning that this generalization retains the greatest advantage of the traditional three-term conjugate gradient method, that is, without relying on the line search, the generated search direction is descending, and the sufficient descent condition (3.3) is satisfied. This serves as a significant supplement to the PRP method, which, as mentioned in Example 3.1, may not independently generate descent directions. Under the common assumptions, the global convergence of the proposed method is obtained by the generalized Wolfe line search procedure. Numerical experiments show that the proposed scheme integrates the advantages of the PRP method and the SD method, i.e., the search direction exhibits both the conjugate property and the sufficient descent property.
Since the extension of the conjugate gradient method to vector optimization may lead to the loss of the descent property, it would be worthwhile to study whether the three-term extensions of other conjugate gradient methods in vector optimization can be developed while retaining the properties of scalar optimization. Additionally, developing a general three-term conjugate gradient method for vector optimization remains an open problem.
Declarations
The author declares to have no competing interests.